OpenStudy (loser66):

Let S is a subset of [0,1] consist of all infinite decimal expansions x =0.a1a2a3...... where all but finitely many digits are 5 or 6. Find sup S Please, help

OpenStudy (loser66):

I know sup S = 1, but don't know how to put it in logic

OpenStudy (anonymous):

me neither but perhaps you could make a sequence like \[.95656...\\ .995656...\\ .9995656...\] which follows the prescription and has a sup of one

OpenStudy (loser66):

\([0,1]\supset S\) and 1 is sup ([0,1]), so that 1 is upper bound of S Let \(\varepsilon >0\) \(x>1-\varepsilon\) \(x = 0.999999.....(n times) ..5 > 1-10^n\)

OpenStudy (loser66):

@kirbykirby

OpenStudy (kirbykirby):

I'm kind of in the same boat as @satellite73 on this one. Since you can get as close to as 1 as you want with those sequences, then it makes sense for \(\sup S =1\), but I'm not too sure how to formalize this one :\ I guess it would be somewhat analogous to a situation where you have \(\sup\{ x \in \mathbb{R}\mid 0 < x < 1 \}=1\) But yeah, technically, you could specify some epsilon the way you did (though I think you mean \(10^{-n}\) ;) ), and you can always find some \(\varepsilon \) in that way as there exists \(\alpha \in S\) such that \(\alpha > 1-\varepsilon = 1-10^{-n} \) for any \(n\) as large as you want it to be, and \(\varepsilon\) will still be \(> 0\)

OpenStudy (ikram002p):

its okey to have sup of 1